OpenStudy (anonymous):

what does this mean ??? "write in the form y=mx+b and state the slope and y-intercept: 7x-6y+5=0

OpenStudy (anonymous):

isolate y m = slope b = y-intercept

OpenStudy (anonymous):

ok.. I will attempt to do so. Do you think this question requires a graph?

OpenStudy (anonymous):

No graph. Rewrite it by keeping y on the left side and the rest on the right side.

OpenStudy (anonymous):

ok, Thankyou! So to start my question will automatically look like this -6y=-7x+5?? Does this mean than that 7 is my slope?? or do I need to divide everything b y -6

OpenStudy (phi):

Start with 7x-6y+5=0 add 6y to both sides of the equation then divide both sides of the equation by 6 (remember when dividing, you divide every term by 6)

OpenStudy (radar):

Miss_Q, you are on the right track, except for the sign of the "5" and yes you will divide all terms by -6.

OpenStudy (anonymous):

when dividing all terms wouldnt it by +6?? I have done this and now my answer looks like this 1.16x+.83=y....

OpenStudy (phi):

Generally, people leave the answers as fractions, unless you need decimals.

OpenStudy (radar):

Let me rearrange that y=1.16x+.833 or leaving as fractions y=7/6(x) +5/6 That looks to be correct.

OpenStudy (phi):

You now know m and b. The m is the number (fraction) in front of the x

OpenStudy (radar):

m is the coefficient of x b is the constant m= 7/6 b=5/6

OpenStudy (anonymous):

So now if I were to state the slope it would be (m) and b represents the y intercept!! I really appreciate your help radar and phi!

OpenStudy (anonymous):

I think I over think it... always in fraction form!

OpenStudy (radar):

Your teacher may prefer fractional form.